The length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.
Answer by prabhjyot(164)

Let the width of the field be x ft.
So length = (2x-5) ft
Perimeter = 2[ length + width]
2[2x-5+x] = 230
2{3x-5} = 230
6x - 10 = 230
6x = 240
x = 40
Thus the width of the field = 40 ft.
Length = 2x-5
= 2*40 - 5
= 80 - 5
= 75
Width = 40 ft and length = 75 ft